# Verifiable Natural Language to Linear Temporal Logic Translation: A Benchmark Dataset and Evaluation Suite

### v1

@@ A.4 Grounding Prompts
Figure 3: Few-shot examples for the Warehouse Scenario
Figure 4: Few-shot examples for the Traffic Light Scenario
Figure 5: Few-shot examples for the Search and Rescue Scenario
A.5 Scenario Configurations
@@
A.6 Compute Resources and External Code and License Information
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims in the beginning of the paper are justified by the experiments and details provided in the main body of the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitations are discussed in Section 5.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: We do not provide any theories. All results are experimental, and all experiments are provided.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: The paper is thorough in its description of the experiments and the data. Someone familiar with the area could feasibly replicate this work given the citations to existing work and the details provided. Regardless, we provide all code and the dataset is public and provided as required.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: The dataset is public and the code and experimental details will be in the supplemental material.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Full details are in the code, but the information in the paper is sufficient for understanding the work and the meaning of the evaluations.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: Our evaluations present only the mean of each metric. Our evaluations are (relatively) small, composed of <1000 examples, and we focus on the presentation and descriptions of our new datasets, rather than a detailed review of current frameworks.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: See Appendix A.6.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We have no ethical issues and have read the code.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss the benefits to NL to TL translation which can be used for improved safe operation of formally specified systems.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A] .
  44. Justification: Our dataset does not enable dangerous models or model behavior.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: See Appendix A.6.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes] .
  52. Justification: Dataset is published as required. All code released in the supplemental material is well documented.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: No human subjects
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No human subjects
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Justification: All uses of LLM for help with dataset generation and for experimental evaluation of methods which use LLMs were disclosed. In addition we have provided the prompts for our unique usage.

### v2

@@ A.6 Grounding Prompts
Few-shot Prompt:
”role”: ”system”, ”content”: ”You are an LTL translation assistant, your goal is to return the desired prop_dict, a dictionary that relates natural language atomic proposition/predicate references to their canonical/known representation in the scenario.”, ”role”: ”user”, ”content”: Few-shot Examples: {examples from ALL domains, shown in appendix A.7, total of 9 examples} Now predict: Sentence: {sentence} Lifted: {lifted_sentence} Prop_dict:
Scenario Prompt:
”role”: ”system”, ”content”: ”You are an LTL translation assistant, your goal is to return the desired prop_dict, a dictionary that relates natural language atomic proposition/predicate references to their canonical/known representation in the scenario.”, ”role”: ”user”, ”content”: Scenario Configuration: scenario yaml, given in appendix A.8 Few-shot Examples: {examples from this specific scenario, shown in Appendix A.7} Now predict: Sentence: {sentence} Lifted: {lifted_sentence} Prop_dict:
A.7 Few-shot Examples by Scenario
Warehouse Examples
Sentence: [”The system must eventually, avoid prop_1”] Lifted Sentence: [”The system must eventually, avoid prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”deliver”, ”action_ref”: ”drop off”, ”args_canon”: [”sandwich loading_dock”], ”args_ref”: [”square food loading dock”] } } Sentence: [”Whenever prop_1 holds, prop_2 holds as well.”] Lifted Sentence: [”Whenever prop_1 holds, prop_2 holds as well.”] prop_dict: { ”prop_1”: { ”action_canon”: ”idle”, ”action_ref”: ”remain still”, ”args_canon”: [], ”args_ref”: [] }, ”prop_2”: { ”action_canon”: ”get_help”, ”action_ref”: ”call for help”, ”args_canon”: [], ”args_ref”: [] } } Sentence: [”If prop_2 holds, then in the next step prop_3 persists until prop_1 holds, or else prop_3 holds forever.”] Lifted Sentence: [”If prop_2 holds, then in the next step prop_3 persists until prop_1 holds, or else prop_3 holds forever.”] prop_dict: { ”prop_1”: { ”action_canon”: ”pickup”, ”action_ref”: ”grab”, ”args_canon”: [”hot_dog”], ”args_ref”: [”bunned sausage”] }, ”prop_2”: { ”action_canon”: ”pickup”, ”action_ref”: ”grab”, ”args_canon”: [”potted_plant”], ”args_ref”: [”plant”] }, ”prop_3”: { ”action_canon”: ”search”, ”action_ref”: ”search for”, ”args_canon”: [”cup”], ”args_ref”: [”beverage cup”] } } Search and Rescue Examples
Sentence: [”This controller must always avoid prop_1”] Lifted Sentence: [”This controller must always avoid prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”record”, ”action_ref”: ”begin recording”, ”args_canon”: [”fire_source”], ”args_ref”: [”fire source”] } } Sentence: [”In this task, take a photo of flood, then return home.”] Lifted Sentence: [”In this task, prop_1 then prop_2”] prop_dict: { ”prop_1”: { ”action_canon”: ”photo”, ”action_ref”: ”take a photo of”, ”args_canon”: [”flood”], ”args_ref”: [”flood”] }, ”prop_2”: { ”action_canon”: ”go_home”, ”action_ref”: ”return home”, ”args_canon”: [], ”args_ref”: [] } } Sentence: [”If every record flood is eventually followed by talking to the safe victim, then avoid the impending debris must occur infinitely often.”] Lifted Sentence: [”If every prop_1 is eventually followed by prop_2 then prop_3 must occur infinitely often.”] prop_dict: { ”prop_1”: { ”action_canon”: ”record”, ”action_ref”: ”record”, ”args_canon”: [”flood”], ”args_ref”: [”flood”] }, ”prop_2”: { ”action_canon”: ”communicate”, ”action_ref”: ”talk to”, ”args_canon”: [”safe_victim”], ”args_ref”: [”safe victim”] }, ”prop_3”: { ”action_canon”: ”avoid”, ”action_ref”: ”avoid”, ”args_canon”: [”impending_debris”], ”args_ref”: [”impending debris”] } }
Traffic Light Examples
Sentence: [”You”, ”must”, ”eventually,”, ”avoid”, ”set”, ”east”, ”light”, ”yellow.”] Grounded: [”You”, ”must”, ”eventually,”, ”avoid”, ”prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”change”, ”action_ref”: ”set”, ”args_canon”: [”light_east”, ”yellow”], ”args_ref”: [”east light”, ”yellow”] } } Sentence: [”Both”, ”change”, ”west”, ”light”, ”red”, ”and”, ”take”, ”a”, ”video”, ”of”, ”the”, ”car”, ”on”, ”southwest”, ”10th”, ”avenue”, ”hold”, ”at”, ”every”, ”step.”] Grounded: [”Both”, ”prop_1”, ”and”, ”prop_2”, ”hold”, ”at”, ”every”, ”step.”] prop_dict: { ”prop_1”: { ”action_canon”: ”change”, ”action_ref”: ”change”, ”args_canon”: [”light_west”, ”red”], ”args_ref”: [”west light”, ”red”] }, ”prop_2”: { ”action_canon”: ”record”, ”action_ref”: ”take a video of”, ”args_canon”: [”car”, ”southwest_10th_avenue”], ”args_ref”: [”car”, ”southwest 10th avenue”] } } Sentence: [”If”, ”take”, ”a”, ”picture”, ”of”, ”the”, ”car”, ”on”, ”northwest”, ”6th”, ”street”, ”holds”, ”and”, ”set”, ”east”, ”light”, ”green”, ”holds”, ”next,”, ”then”, ”request”, ”assistance”, ”holds”, ”in”, ”the”, ”step”, ”after”, ”that.”] Grounded: [”If”, ”prop_1”, ”holds”, ”and”, ”prop_2”, ”holds”, ”next,”, ”then”, ”prop_3”, ”holds”, ”in”, ”the”, ”step”, ”after”, ”that.”] prop_dict: { ”prop_1”: { ”action_canon”: ”photo”, ”action_ref”: ”take a picture of”, ”args_canon”: [”car”, ”northwest_6th_street”], ”args_ref”: [”car”, ”northwest 6th street”] }, ”prop_2”: { ”action_canon”: ”change”, ”action_ref”: ”set”, ”args_canon”: [”light_east”, ”green”], ”args_ref”: [”east light”, ”green”] }, ”prop_3”: { ”action_canon”: ”get_help”, ”action_ref”: ”request assistance”, ”args_canon”: [], ”args_ref”: [] } }
A.8 Scenario Configurations
@@
Figure 6: Kitchen Assistant Scenario Configuration file
A.9 Compute Resources and External Code and License Information
@@
A.10 Large Language Model Disclosure
During the preparation of this paper, the authors employed large language models (LLMs) as assistive tools for limited tasks including proof-reading, text summarization, and the discovery of related work. All substantive research contributions, analyses, and claims presented in this paper were conceived, developed, and verified by the authors. The authors maintain full ownership and responsibility for the content of the paper, including its technical correctness, originality, and scholarly contributions.